# Introducing Routing Functions to Vision-Language Parameter-Efficient Fine-Tuning with Low-Rank Bottlenecks

### v1

@@ 4.3 Results
I-Tuning ituning2023 implements an Adapter with cross-attention for image captioning. To have a fair comparison, we follow the settings of I-Tuning for both LoRA and Adapter. Specifically, (1) instead of using the visual features, we prepend the "<s>" token to the input text. (2) We use the last hidden state from ViT as key and value for cross-attention and as $x_{R}$ for routing functions. This change is needed, since the [CLS] feature is of length 1, which does not work for cross-attention due to the softmax function.
In the case of LoRA, we also implement a variant with separate $W_{down}$ mappings for $x_{R}$ and $x_{H}$. This brings LoRA with routing function closer to LoRA with cross-attention.999 Appendix 0.A shows detailed architectures. We down-project to $r/4$ and $r$ for LoRA w/ & w/o separate mapping, respectively, to have the same level of parameter count. Without separate mapping refers to the setting used in previous experiments, where $x_{R}$ and $x_{H}$ share the same $W_{down}$. As shown in Tab. 5, despite of having less parameters, LoRA+$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ (w/ separate mapping) achieves significantly better performance than LoRA with cross-attention. The setup of LoRA+$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ (w/o separate mapping) outperforms LoRA with cross-attention by a large margin using a similar amount of parameters.
For Adapter, the cross-attention module has the same intermediate dimension $r$ as the $W_{down}$ in Adapter with routing function. Naturally, the separate query/key/value mappings make that the Adapter with cross-attention has more than double the number of parameters compared to the Adapter equipped with the routing function ($x_{t}(x_{v})^{\mathsf{T}}x_{v}$). Tab. 5 shows that both achieve a comparable performance.
@@
Incorporating nonlinear activation functions into routing functions We have proposed linear routing functions for efficiency reasons, but they do not need to be exclusively linear. Taking inspirations from the linear attention variants pmlr-v119-katharopoulos20a ; choromanski2021rethinking ; zhen2022cosformer , where the linear operation is often equipped with nonlinear activation functions to ensure non-negativity for similarity scores between query and key, we explore whether a similar operation affects the representation power of the routing functions in the low-rank bottleneck. For routing function $x_{t}(x_{v})^{\mathsf{T}}x_{v}$, we use the nonlinear activation function $ReLU(\cdot)$ as in previous work zhen2022cosformer , and transform it to $ReLU(x_{t})ReLU(x_{v})^{\mathsf{T}}x_{v}$. The results are presented in Tab. 6. With or without $ReLU(\cdot)$, our routing function achieves similar performance on VQA and image captioning.
Qualitative Analysis We first qualitatively compare the VL alignment pattern of LoRA-tuned GPT2 with or without a routing function on image captioning. An example is presented in Figure 4. With the addition of routing function $x_{t}(x_{v})^{\mathsf{T}}x_{v}$, the model pays more attention to the nouns (e.g. man, a big screen) and the verbs (e.g. watching), which are essential in learning the VL alignment. At the same time, it pays less attention to meaningless words like "something". This is in line with the finding that the routing function stresses components of $x_{t}$ that are aligned in the image.
@@
Next we qualitatively examine some examples from the inference phase for VQA and Image Captioning. We present examples in Figure 5 where conventional LoRA fails. For VQA, previous experiments in Tab. 3 indicate that our routing functions work particularly well on color related questions. As also shown in Figure 5, conventional LoRA fails to align the color red to the lollipop. The model tuned with conventional LoRA also fails to correctly answer other types of questions as presented in the figure. Without routing functions, captions generated from the LM tuned with conventional LoRA often make mistakes in comprehending the visual information. Occasionally, we observe a sequence of blank spaces as presented in the first example using conventional LoRA.
@@ 5 Experiments with Encoder-Decoder Language Models
In this section, we study the VL PEFT tasks in a more complicated setting, where we consider an encoder-decoder LM with an additional mapping network dedicated for the vision encoder. The model is trained in a multi-task learning setting, where four different VL tasks are jointly learned, as in cho2021vlt5 ; Sung_2022_CVPR . Despite the increased complexity of optimizing the network in this scenario, we show that integrating routing functions into the low-rank bottlenecks yields performance enhancements. Moreover, consistent patterns are observed for different scenarios, which shed lights on future research to improve VL PEFT tasks.
@@ 5.1 Experimental Design
Following Sung_2022_CVPR , we use the CLIP-BART architecture to study the effect of our routing functions on the VL PEFT tasks in a multitask learning setting. CLIP-BART is a widely used architecture for VL tasks. We present the simplified model architecture of CLIP-BART with Adapters and the integrated routing functions in Figure 6. The original CLIP-BART uses a trainable visual projection module to map the visual features extracted by CLIP to the embedding space of BART lewis-etal-2020-bart . CLIP-BART uses task-specific prompts as the textual input: For example, in Figure 6, "<vqa>" is added as prompt to the text for VQA. The mapped visual features are then prepended to the textual input. Together, they serve as the input to BART. CLIP-BART is trained and evaluated in a multitask setting, that is, VQAv2 balanced_vqa_v2 and GQA Hudson_2019_CVPR for visual question answering, $\text{NLVR}^{2}$ suhr-etal-2019-corpus using natural language for visual reasoning, and COCO Cap. mscoco for image captioning. CLIP-BART unifies the four tasks into one text generation problem following cho2021vlt5 , where four different task prompts distinguish the task types. CLIP-BART can be considered as the representative of the current state-of-the-art on VL PEFT tasks.
Figure 6: Architecture of CLIP-BART with Adapter. The standard Adapter architecture as shown in Footnote 2 is adopted. Only the modules in green are trainable. To integrate routing functions into the model, we use the pooled last encoder hidden state as feature $x_{R}$, which is added to the Adapters in the decoder.
To integrate our routing functions into the model, two types of inputs $x_{R}$ can be considered, namely the visual features after the visual projection, and the last encoder hidden state as presented in Figure 6. The visual features only contain visual information. Because the last encoder hidden state contains both textual and visual information, it can only be added to the decoder. We experiment with both LoRA and Adapter as the PEFT method for fine-tuning CLIP-BART. In this multitask learning setting, one can use (1) a single LoRA/Adapter shared across tasks or (2) multiple LoRAs/Adapters without parameter sharing, where each task has its own LoRA/Adapter. We start by exploring the impact of different routing functions on the single LoRA/Adapter setting. Then, based on the observations obtained in the setting with single Adapter, we explore ways of using the routing functions in the multiple Adapters setting.
@@ 5.2 Implementation Details
Following Sung_2022_CVPR , we use CLIP-ResNet101 as our vision encoder pmlr-v139-radford21a . The same procedure for extracting visual features is adopted as in Sung_2022_CVPR . We use $\text{BART}_{\text{base}}$ as our encoder-decoder LM as in Sung_2022_CVPR . Following Sung_2022_CVPR , we use $r=96$ for all experiments. The training pipeline in Sung_2022_CVPR is used for a fair comparison. Specifically, (1) we use the same seed for all experiments. (2) Due to limited resources, we do not perform an extensive hyper-parameter search as in Sung_2022_CVPR . (3) We train the model with AdamW optimizer and linear decay for 20 epochs. We warm-up the learning rate from 0 to the highest in the first 2 epochs. (4) The batch size is set to 500. In cases where we compare results with Hu_2023_ICCV ; Sung_2022_CVPR , we use the same seeds as in their publication. For more details, we refer to Appendix 0.A.
@@ 5.3 Results
Functions VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg. VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg.
$\text{None}^{\dagger}$ 65.15 53.66 72.58 115.01 76.60 65.76 54.16 73.19 114.61 76.93
$\text{None}^{\ddagger}$ 65.2 53.6 71.9 115.3 76.5 65.9 54.5 74.2 114.9 77.4
$x_{t}\circ x^{\prime}_{v}$ $\mathbf{65.68}$ $\mathbf{53.90}$ $\mathbf{73.96}$ 113.63 76.79 65.70 $\mathbf{54.18}$ $\mathbf{74.47}$ 114.02 77.09
$x_{t}+x^{\prime}_{v}$ $\underline{65.23}$ $\underline{53.45}$ 73.25 115.54 76.87 $\underline{65.81}$ 53.63 73.99 113.69 76.78
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 65.00 53.21 $\underline{73.88}$ $\underline{117.43}$ $\underline{77.38}$ 65.77 $\underline{54.14}$ $\underline{74.29}$ $\underline{117.99}$ $\mathbf{78.05}$
$x_{t}x^{\prime\prime}_{v}$ 64.84 52.72 73.46 $\mathbf{119.53}$ $\mathbf{77.64}$ $\mathbf{65.82}$ 53.37 72.33 $\mathbf{118.94}$ $\underline{77.62}$
@@
Experiments with multiple Adapters We further explore the behavior of the routing functions in models with multiple Adapters, where we train a separate Adapter per task. As also observed in Sung_2022_CVPR , using multiple Adapters increases the complexity of the model, and yields worse performance compared to using a single Adapter. As presented in Tab. 8, using the four routing functions with multiple Adapters leads to comparable or better performance as compared to multiple Adapters with no routing functions. Moreover, in line with what we discovered in the experiments with models with a single Adapter, different routing functions shine on different types of tasks.
In this set-up, we also experiment with combinations of routing functions, which are determined based on the results obtained with a single Adapter. The four combinations of routing functions integrate the routing functions that work best for each individual task (e.g. $x_{t}x^{\prime}_{v}$ for COCO Cap.). The results in Tab. 8 evidence that in this multitask setting combinations of routing functions that each work well for an individual task yield an improved performance compared to using no routing functions.
@@
Functions VQA GQA $\text{NLVR}^{2}$ COCO Cap. VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg.
$\text{None}^{\ddagger}$ 65.4 54.0 69.8 114.3 75.9
$x_{t}\circ x^{\prime}_{v}$ 65.61 53.51 70.78 113.72 75.91
$x_{t}+x^{\prime}_{v}$ 65.19 $\underline{53.84}$ $\underline{70.99}$ 115.64 76.42
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 64.93 52.11 70.25 116.20 75.87
$x_{t}x^{\prime\prime}_{v}$ 65.32 52.67 70.58 $\underline{117.92}$ 76.62
comb1 $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $x_{t}x^{\prime\prime}_{v}$ $\mathbf{66.37}$ $\mathbf{54.21}$ 69.27 116.94 76.70
comb2 $x_{t}+x^{\prime}_{v}$ $x_{t}+x^{\prime}_{v}$ $x_{t}(x_{v})^{T}x_{v}$ $x_{t}x^{\prime\prime}_{v}$ 65.14 53.37 70.88 117.58 76.74
comb3 $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}x^{\prime\prime}_{v}$ $\underline{66.34}$ 53.77 $70.02$ 117.73 $\mathbf{76.97}$
comb4 $x_{t}+x^{\prime}_{v}$ $x_{t}+x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}x^{\prime\prime}_{v}$ 65.47 52.99 $\mathbf{71.01}$ $\mathbf{118.34}$ $\underline{76.95}$
@@ 6 Conclusion
In this work, we have introduced routing functions as helper functions to guide the feature learning in low-rank bottlenecks in architectures that are typical for parameter-efficient fine-tuning (PEFT) such as LoRA and Adapter. We have shown their validity in vision-language tasks, where we have focused on linear functions that do not require any extra parameters. We have demonstrated performance gains in single-task and multi-task settings.
@@ Appendix 0.A Additional Implementation Details
For experiments on VQAv2 balanced_vqa_v2 in Sec. 4.3(Tab. 3), we list the specific question types involved in "what", "color", "time" and "person" below:
@@
Architectures with separate mapping We present the illustration of the architectures used for experiments on comparing the routing functions to cross-attention in Figure A7. The experiments can be found in Sec. 4.3 (Tab. 5). We also implement a variant with separate $W_{down}$ mappings for $x_{R}$ and $x_{H}$.
To keep the number of parameters comparable to our standard setting, in LoRA + cross-attention as shown in Figure A7 (c), we down-project to $r/4$. After cross-attention, we first up-project to rank $r$, which is the same as the standard setting. Finally we up project back to the original dimension $d$. In that case, to have a fair comparison, in LoRA + routing functions (w/ separate mapping) as shown in Figure A7 (d), we follow exactly the same mapping strategy, namely $d\rightarrow r/4\rightarrow r\rightarrow d$, which results in less trainable parameters as compared to our standard setting and the cross-attention in Figure A7 (c). For detailed parameter count, we refer to Tab. 5 in the main text.
Figure A7: Illustration of architectures used in comparison to cross-attention. (a) Standard way of integrating routing functions in PEFT methods. We implement this in Adapter/LoRA w/o separate mapping. (b) Cross-attention integrated in Adapter as introduced in ituning2023 . We implement this in Adapter + cross-attention following ituning2023 . (c) Cross-attention in PEFT, but with additional up-projection map which project the hidden states to $r$ first. We implement this in LoRA with cross-attention. (d) PEFT with routing funcitons using separate mapping for vision and language. We implement this in LoRA with separate mapping to examine how will separate mappings in cross-attention affect performance.
Additional implementation details for CLIP-BART In addition to Sec. 5.2, we report additional implementation details of CLIP-BART in this part. Following Sung_2022_CVPR , we extract the 7$\times$7 grid features from the last convolutional layer of the vision encoder, which we down-project to 6$\times$6 use adaptive maximum-pooling. The image resolution is set to 224$\times$224 following Sung_2022_CVPR . For CLIP-BART experiments, we conduct our experiments based on the official implementation of VL-Adapter Sung_2022_CVPR , which searches for optimal learning rate among $1\times 10^{-4}$, $3\times 10^{-4}$, $1\times 10^{-3}$}. Limited to resources, we only search for optimal learning rate among $3\times 10^{-4}$, $1\times 10^{-3}$}. The adopted learning rates for different experiments are presented in Tab. A9.

### v2

@@ 4.3 Results
I-Tuning [27] implements an Adapter with cross-attention for image captioning. To have a fair comparison, we follow the settings of I-Tuning for both LoRA and Adapter. Specifically, (1) instead of using the visual features, we prepend the "<s>" token to the input text. (2) We use the last hidden state from ViT as key and value for cross-attention and as $x_{R}$ for routing functions. This change is needed, since the [CLS] feature is of length 1, which does not work for cross-attention due to the Softmax function.
In the case of LoRA, we also implement a variant with separate $W_{down}$ mappings for $x_{R}$ and $x_{H}$. This brings LoRA with routing function closer to LoRA with cross-attention.999 Appendix 0.A shows detailed architectures and explains why routing functions outperform cross-attention. We down-project to $r/4$ and $r$ for LoRA w/ & w/o separate mapping, respectively, to have the same level of parameter count. Without separate mapping refers to the setting used in previous experiments, where $x_{R}$ and $x_{H}$ share the same $W_{down}$. As shown in Tab. 5, despite of having less parameters, LoRA+$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ (w/ separate mapping) achieves significantly better performance than LoRA with cross-attention. The setup of LoRA+$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ (w/o separate mapping) outperforms LoRA with cross-attention by a large margin using a similar amount of parameters.
For Adapter, the cross-attention module has the same intermediate dimension as the $W_{down}$ in Adapter with routing function. Naturally, the separate query/key/value mappings make that the Adapter with cross-attention has more than double the number of parameters compared to the Adapter equipped with the routing function ($x_{t}(x_{v})^{\mathsf{T}}x_{v}$). Both achieve comparable performance (Tab. 5).
@@
Incorporating nonlinear activation functions into routing functions We have proposed linear routing functions for efficiency reasons, but they do not need to be exclusively linear. Taking inspirations from the linear attention variants [15, 4, 30], where the linear operation is often equipped with nonlinear activation functions to ensure non-negativity for similarity scores between query and key, we explore whether a similar operation affects the representation power of the routing functions in the low-rank bottleneck. For routing function $x_{t}(x_{v})^{\mathsf{T}}x_{v}$, we use the nonlinear activation function $ReLU(\cdot)$ as in previous work [30], and transform it to $ReLU(x_{t})ReLU(x_{v})^{\mathsf{T}}x_{v}$. The results are presented in Tab. 6. With or without $ReLU(\cdot)$, our routing function achieves similar performance on VQA and image captioning.
Qualitative Analysis We first qualitatively compare the VL alignment pattern of LoRA-tuned GPT2 with or without a routing function on image captioning. An example is presented in Fig. 4. With the addition of routing function $x_{t}(x_{v})^{\mathsf{T}}x_{v}$, the model pays more attention to the nouns (e.g. man, a big screen) and the verbs (e.g. watching), which are essential in learning the VL alignment. At the same time, it pays less attention to meaningless words like "something". This is in line with the finding that the routing function stresses components of $x_{t}$ that are aligned in the image.
@@
Next we qualitatively examine some examples from the inference phase for VQA and Image Captioning. We present examples in Fig. 5 where conventional LoRA fails. For VQA, previous experiments in Tab. 3 indicate that our routing functions work particularly well on color related questions. As also shown in Fig. 5, conventional LoRA fails to align the color red to the lollipop. The model tuned with conventional LoRA also fails to correctly answer other types of questions as presented in the figure. Without routing functions, captions generated from the LM tuned with conventional LoRA often make mistakes in comprehending the visual information. Occasionally, we observe a sequence of blank spaces as presented in the first example using conventional LoRA.
@@ 5 Experiments with Encoder-Decoder Language Models
In this section, we study the VL PEFT tasks in a more complicated setting, where we consider an encoder-decoder LM with an additional mapping network dedicated for the vision encoder. The model is trained in a multi-task learning setting, where four different VL tasks are jointly learned, as in [3, 37]. Despite the increased complexity of optimizing the network in this scenario, we show that integrating routing functions into the low-rank bottlenecks yields performance enhancements. Moreover, consistent patterns are observed for different scenarios, which shed lights on future research to improve VL PEFT tasks.
@@ 5.1 Experimental Design
Following [37], we use the CLIP-BART architecture to study the effect of our routing functions on the VL PEFT tasks in a multitask learning setting. CLIP-BART is a widely used architecture for VL tasks. We present the simplified model architecture of CLIP-BART with Adapters and the integrated routing functions in Fig. 6. The original CLIP-BART uses a trainable visual projection module to map the visual features extracted by CLIP to the embedding space of BART [20]. CLIP-BART uses task-specific prompts as the textual input: For example, in Fig. 6, "<vqa>" is added as prompt to the text for VQA. The mapped visual features are then prepended to the textual input. Together, they serve as the input to BART. CLIP-BART is trained and evaluated in a multitask setting, that is, VQAv2 [7] and GQA [12] for visual question answering, $\text{NLVR}^{2}$ [35] using natural language for visual reasoning, and COCO Cap. [23] for image captioning.
Figure 6: Architecture of CLIP-BART with Adapter. The standard Adapter architecture is adopted (see Footnote 2). Only the modules in green are trainable. To apply routing functions, we use the pooled last encoder hidden state as feature $x_{R}$, which is added to the Adapters in the decoder.
CLIP-BART unifies the four tasks into one text generation problem following [3], where four different task prompts distinguish the task types. CLIP-BART can be considered as the representative of the current state-of-the-art on VL PEFT tasks.
To integrate our routing functions into the model, two types of inputs $x_{R}$ can be considered, namely the visual features after the visual projection, and the last encoder hidden state as presented in Fig. 6. The visual features only contain visual information. Because the last encoder hidden state contains both textual and visual information, it can only be added to the decoder. We experiment with both LoRA and Adapter as the PEFT method for fine-tuning CLIP-BART. In this multitask learning setting, one can use (1) a single LoRA/Adapter shared across tasks or (2) multiple LoRAs/Adapters without parameter sharing, where each task has its own LoRA/Adapter. We start by exploring the impact of different routing functions on the single LoRA/Adapter setting. Then, based on the observations obtained in the setting with single Adapter, we explore ways of using the routing functions in the multiple Adapters setting.
@@ 5.2 Implementation Details
Following [11, 37], we use CLIP-ResNet101 as our vision encoder [31]. We use $\text{BART}_{\text{base}}$ as our encoder-decoder LM as in [37]. Following [37], we use $r=96$ for all experiments. The training pipeline in [37] is used for a fair comparison. Specifically, (1) we use the same seed for all experiments. (2) Due to limited resources, we do not perform an extensive hyper-parameter search as in [37]. (3) We train the model with AdamW optimizer and linear decay for 20 epochs. We warm-up the learning rate from 0 to the highest in the first 2 epochs. (4) The batch size is set to 500. For more details, we refer to Appendix 0.A.
@@ 5.3 Results
Functions VQA GQA $\text{NLVR}^{2}$ Avg. COCO Cap. VQA GQA $\text{NLVR}^{2}$ Avg. COCO Cap.
$\text{None}^{\dagger}$ 65.15 53.66 72.58 63.80 115.01 65.76 54.16 73.19 64.37 114.61
$x_{t}\circ x^{\prime}_{v}$ $\mathbf{65.68}$ $\mathbf{53.96}$ 73.42 $\mathbf{64.35}$ 113.94 $\mathbf{65.92}$ $\mathbf{54.34}$ $\underline{74.23}$ $\mathbf{64.83}$ 114.38
$x_{t}+x^{\prime}_{v}$ $\underline{65.14}$ $\underline{53.73}$ $\underline{73.51}$ $\underline{64.13}$ 114.96 $\underline{65.89}$ $\underline{54.18}$ 73.90 $\underline{64.66}$ 114.39
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 64.94 53.56 $\mathbf{73.60}$ 64.03 $\underline{117.80}$ 65.84 53.65 $\mathbf{74.31}$ 64.03 $\underline{117.65}$
$x_{t}x^{\prime\prime}_{v}$ 64.84 53.13 72.98 63.65 $\mathbf{119.26}$ 65.83 53.61 73.27 64.24 $\mathbf{118.86}$
@@
Experiments with multiple Adapters We further explore the behavior of the routing functions in models with multiple Adapters, where we train a separate Adapter per task. As also observed in [37], using multiple Adapters increases the complexity of the model, and yields worse performance compared to using a single Adapter. As presented in Tab. 8, using the four routing functions with multiple Adapters leads to comparable or better performance as compared to multiple Adapters with no routing functions. Moreover, in line with what we discovered in the experiments with models with a single Adapter, different routing functions shine on different types of tasks.
In this set-up, we also experiment with combinations of routing functions. The four combinations integrate the routing functions that work best for each individual task (e.g. $x_{t}x^{\prime}_{v}$ for COCO Cap.). The results in Tab. 8 evidence that in this multitask setting combinations of routing functions yield an improved performance compared to using no routing functions. GQA requires more fine-grained image representation. For efficiency reason, we only adopt $x_{v}$ as the down-projection from the pooled last encoder hidden state. Nevertheless, comb1 can recover from this shortcoming. More detailed image representations as used in the cross-attention experiments of Tab. 5 might still enhance the results.
@@
Functions VQA GQA $\text{NLVR}^{2}$ COCO Cap. VQA GQA $\text{NLVR}^{2}$ Avg. COCO Cap.
$\text{None}^{\ddagger}$ 65.4 54.0 69.8 63.07 114.3
$x_{t}\circ x^{\prime}_{v}$ 65.61 53.51 70.78 $\underline{63.30}$ 113.72
$x_{t}+x^{\prime}_{v}$ 65.19 $\underline{53.84}$ $\underline{70.99}$ 63.34 115.64
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 64.93 52.11 70.25 62.43 116.20
$x_{t}x^{\prime\prime}_{v}$ 65.32 52.67 70.58 62.86 $\underline{117.92}$
comb1 $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $x_{t}x^{\prime\prime}_{v}$ $\mathbf{66.37}$ $\mathbf{54.21}$ 69.27 63.28 116.94
comb2 $x_{t}+x^{\prime}_{v}$ $x_{t}+x^{\prime}_{v}$ $x_{t}(x_{v})^{T}x_{v}$ $x_{t}x^{\prime\prime}_{v}$ 65.14 53.37 70.88 63.13 117.58
comb3 $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}x^{\prime\prime}_{v}$ $\underline{66.34}$ 53.77 $70.02$ $\mathbf{63.38}$ 117.73
comb4 $x_{t}+x^{\prime}_{v}$ $x_{t}+x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}x^{\prime\prime}_{v}$ 65.47 52.99 $\mathbf{71.01}$ 63.16 $\mathbf{118.34}$
@@ 6 Conclusion
We have introduced feature routing functions to help guide the feature learning in low-rank bottlenecks in architectures that are typical for parameter-efficient fine-tuning such as LoRA and Adapter. We have shown their validity in vision-language tasks, where we have focused on linear functions that do not require any extra parameters. We have demonstrated performance gains in single-task and multi-task settings in combination with a variety of foundation models.
@@
Acknowledgements
This work is funded by the Flanders AI Research Program, European Research Council Advanced Grant H2020-ERC-2017-ADG 788506, and the China Scholarship Council.
@@ Appendix 0.A Additional Implementation Details
For experiments on VQAv2 [7] in Sec. 4.3(Tab. 3), we list the specific question types involved in "what", "color", "time" and "person" below:
@@
Architectures with separate mapping We present the illustration of the architectures used for experiments on comparing the routing functions to cross-attention in Fig. A7. The experiments can be found in Sec. 4.3 (Tab. 5). We also implement a variant with separate $W_{down}$ mappings for $x_{R}$ and $x_{H}$.
To keep the number of parameters comparable to our standard setting, in LoRA + cross-attention as shown in Fig. A7 (c), we down-project to $r/4$. After cross-attention, we first up-project to rank $r$, which is the same as the standard setting. Finally we up project back to the original dimension $d$. In that case, to have a fair comparison, in LoRA + routing functions (w/ separate mapping) as shown in Fig. A7 (d), we follow exactly the same mapping strategy, namely $d\rightarrow r/4\rightarrow r\rightarrow d$, which results in less trainable parameters as compared to our standard setting and the cross-attention in Fig. A7 (c). For detailed parameter count, we refer to Tab. 5 in the main text.
Figure A7: Illustration of architectures used in comparison to cross-attention. (a) Standard way of integrating routing functions in PEFT methods. We implement this in Adapter/LoRA w/o separate mapping. (b) Cross-attention integrated in Adapter as introduced in [27]. We implement this in Adapter + cross-attention following [27]. (c) Cross-attention in PEFT, but with additional up-projection map which project the hidden states to $r$ first. We implement this in LoRA with cross-attention. (d) PEFT with routing funcitons using separate mapping for vision and language. We implement this in LoRA with separate mapping to examine how will separate mappings in cross-attention affect performance.
Additional implementation details for RoBERTa/GPT2 + ViT In addition to Sec. 4.2, we report additional implementation details of RoBERTa/GPT2 + ViT in this part. We use the Karpathy split of VQAv2 [7] and COCO Captioning [23] (COCO Cap.) in our experiments. For VQAv2, only the parameters of the classifier and the PEFT modules in RoBERTa are updated during training. For COCO Cap., only the parameters of the PEFT modules in GPT2 are updated during training. The maximum length for captions is set to 50. After training, we evaluate the models using the final checkpoint for both tasks.
Additional implementation details for CLIP-BART In addition to Sec. 5.2, we report additional implementation details of CLIP-BART in this part. The same procedure for extracting visual features is adopted as in [37]. Following [37], we extract the 7$\times$7 grid features from the last convolutional layer of the vision encoder, which we down-project to 6$\times$6 use adaptive maximum-pooling. The image resolution is set to 224$\times$224 following [37]. For CLIP-BART experiments, we conduct our experiments based on the official implementation of VL-Adapter [37], which searches for optimal learning rate among $1\times 10^{-4}$, $3\times 10^{-4}$, $1\times 10^{-3}$}. Limited to resources, we only search for optimal learning rate among $3\times 10^{-4}$, $1\times 10^{-3}$}. The adopted learning rates for different experiments are presented in Tab. A9.